OpenStudy (anonymous):

n a word game, you choose a tile from a bag, replace it, and then choose another. If there are 28 vowels and 12 consonants, what is the probability you will choose a consonant and then a vowel?

OpenStudy (anonymous):

This is an independent Probability P(consonant) = 12/40 P(vowel)= 28/40 P(consonant,vowel) = 12/40 * 28/40 = 336/1600

OpenStudy (anonymous):

so the answer is 21/100?

OpenStudy (anonymous):

reduced yes

OpenStudy (anonymous):

Thank youuuuu :)!

OpenStudy (anonymous):

no problem